Given a Greenshields linear density speed-relationship, S = 55.0 - 0.45 D. Flow-speed relationship is Flow=1100S-S^2/g and flow-density relationship is Flow=55D-0.45D^2

1.What are the values for the following:

a) free flow speed __________________________

b) jam density ____________________

c) capacity _______________________
Answer #1

S = 55.0 - 0.45 D

the greenshield model

S=S_f-\frac{S_f}{Dj}D

thus

free flow speed (Sf): 55

jam density (Dj)=55/0.45=122

To determine the capacity of the site need to determine the maximum flow:

\frac{d q}{dD}=55-0.9D=0

D=\frac{55}{0.9}=61.1

density at maximum flow 61

Flow=55D-0.45D^2

=55(61.1)-0.45(61.1)^2=1680.5

capacity=1680.5